Find the resultant using the component method for the given vectors
A = 32degrees 3200lbs
B = 36degrees 1300lbs
C = 35degrees 2100lbs
I added the 3 vectors using the head to tail method but not sure how to add three vectors, this is what I did:
Ax =3200lb(cos32 degrees)
=Ax = 2560
Ay = 3200(sin 32 degrees)
=Ay = 1600
Bx = 1300lb (cos 36 degrees)
= Bx = 1051.7
By - 1300 (sin 36 degrees)
Its not finished but is that the way to do it?

Answer by scott8148(6628)

yes, you are doing it correctly ... find the x and y components of the three vectors and then add them

Rx=Ax+Bx+Cx and Ry=Ay+By+Cy ... you can calculate the angle for R using Rx and Ry ... tan=Ry/Rx